OpenStudy (maheshmeghwal9):

If \[x+\frac{1}{x}=2 \cos \alpha.\]Then Prove that;\[x^n+\frac{1}{x^n}=2 \cos n \alpha.\]

OpenStudy (maheshmeghwal9):

Let \[x=e^{i \alpha}\]&\[\frac{1}{x}=e^{-i \alpha}.\]Then \[[x^n+\frac{1}{x^n}]=[e^{i n \alpha}+ e^{-i n \alpha}]=?\] Now wt to do?????

OpenStudy (anonymous):

Looks like De Moivre to me.

OpenStudy (maheshmeghwal9):

So have I to expand it????????????????

OpenStudy (anonymous):

\[ e^{i n \alpha } = (e^{i\alpha} )^n = \text{cis}n\alpha \]

OpenStudy (anonymous):

I haven't tried yes @maheshmeghwal9 , but that's what I would try yes.

OpenStudy (maheshmeghwal9):

oh ok i gt it thanx a lot:)

OpenStudy (anonymous):

you are welcome